# Jeffery Orbits with Noise Revisited

### v1

@@ Abstract
The behavior of non-spherical particles in a shear-flow is of significant practical and theoretical interest. These systems have been the object of numerous investigations since the pioneering work of Jeffery a century ago. His eponymous orbits describe the deterministic motion of an isolated, rod-like particle in a shear flow. Subsequently, the effect of adding noise was investigated. The theory has been applied to colloidal particles, macromolecules, anisometric granular particles and most recently to microswimmers, for example bacteria. We study the Jeffery orbits of elongated particles subject to noise using Langevin simulations and a Fokker-Planck equation. We extend the analytical solution for infinitely thin needles ($\beta=1$) obtained by Doi and Edwards to particles with arbitrary shape factor ($0\leq\beta\leq 1$) and validate the theory by comparing it with simulations. We examine the rotation of the particle around the vorticity axis and study the orientational order matrix. We use the latter to obtain scalar order parameters $s$ and $r$ describing nematic ordering and biaxiality from the orientational distribution function. The value of $s$ (nematic ordering) increases monotonically with increasing Péclet number, while $r$ (measure of biaxiality) displays a maximum value. From perturbation theory we obtain simple expressions that provide accurate descriptions at low noise (or large Péclet numbers). We also examine the orientational distribution in the v-grad v plane and in the perpendicular direction. Finally we present the solution of the Fokker-Planck equation for a strictly two-dimensional (2D) system. For the same noise amplitude the average rotation speed of the particle in 3D is larger than in 2D.
@@ I Introduction
Even a century after its publication, Jeffery’s work is still relevant to many current research fields for a deep understanding of the behavior of sheared elongated particles, including rheology of granular flows and suspensions [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], nanoparticles, e.g. silica micro-rods in channels [15, 16, 17], red blood cells [18], as well as self-assembly and systems of active particles such as microswimmers [19, 20, 21, 22, 23, 24, 22, 23, 25, 26, 27].
@@ II Jeffery Orbit
where ${\bf p}$ is a unit vector specifying the particle orientation, ${\bf E}=(\nabla{\bf u}+\nabla{\bf u}^{T})/2$, ${\bf W}=(\nabla{\bf u}-\nabla{\bf u}^{T})/2$ are the strain rate and vorticity tensors and $\beta$ is the shape factor (Bretherton parameter [28])
@@
for a particle of elongation $e$. See Fig. 1. Note that $\beta=1$ and $\beta=0$ correspond to infinitely elongated and spherical particles respectively, while $\beta=-1$ describes an infinitely thin disk. We can get particular realizations of the Jeffery equations by choosing the flow geometry. For shear in the $xy$ plane, i.e., ${\bf u}=\dot{\gamma}y\hat{\bf x}$, substituting in (1) and using standard spherical polar coordinates ($\theta,\;\phi$) we obtain
Figure 1: Coordinate system used to describe the motion of an ellipsoid of revolution with shear in the $xy$ (v-grad v) plane. The $x,y,z$ axes correspond to the flow, gradient and vorticity directions, respectively.
@@
$e\tan\phi=\tan\left(\frac{\dot{\gamma}t}{e+1/e}+\kappa\right),$ (5)
@@ III Langevin Description
To simulate the Jeffery orbits with noise, we simply combine these equations with those for the Jeffery orbits (the Itô drift being unchanged). For shear in the $xy$ plane we obtain
@@ IV Fokker-Planck description
where ${\bf w}(\theta,\phi)=(0,\dot{\theta},\dot{\phi}\sin\theta)$, $D$ is the rotational diffusion coefficient, assumed constant [41, 43].
It is convenient to solve the Fokker-Planck equation with shear (v-grad v) in the $xz$ plane, which exploits the higher symmetry of the system with spherical coordinates (the reflection symmetry in the $xz$ plane allows one to use real spherical harmonics). We obtain
@@
$\hat{\Lambda}=-\frac{3}{2}\beta\sin(2\theta\cos\phi)\;+\left(\frac{1-\beta}{2}+\beta\cos^{2}\theta\right)\cos\phi\frac{\partial}{\partial\theta}-\frac{1+\beta}{2}\cot\theta\sin\phi\frac{\partial}{\partial\phi}$ (21)
@@
$\hat{\Lambda}=\beta\hat{\Lambda}_{1}+\frac{1-\beta}{2}i\hat{L}_{y}$ (24)
For $\beta=0$, corresponding to a sphere, Eq. (20) simplifies to
@@
$Dl(l+1)b_{lm}=-\sum_{l^{\prime}\geq 0,\;{\rm even}}^{lmax}\sum_{m^{\prime}=0}^{l^{\prime}}(lm|\hat{\Lambda}|l^{\prime}m^{\prime})b_{l^{\prime}m^{\prime}}$ (30)
From this we generate $l_{max}(1+l_{max}/4)$ simultaneous equations that can be solved to find the coefficients $\{b_{lm}\}$.
@@ V.2 Rotation around the vorticity axis
We first examine the rotation of the particle around the vorticity axis. For shear in the $xy$ plane, in the steady state we have $\langle d\theta/dt\rangle=0$ and
@@
With zero noise we find using Eq. (5)
@@
Interestingly this implies that the infinitely thin particles, $\beta=1$, do not rotate at all, while a spherical particle, $\beta=0$, rotates at $\dot{\gamma}/2$. Fig. 3 shows some simulation results and, for all values of $\beta$, the rate of rotation decreases with increasing noise. Perturbation theory, as developed in Appendix C, can be used to calculate the rotation speed in the limit of large noise (or small Pe).
@@ V.3 Orientational distribution function
Figure 4: Density plots of the particle orientation from Langevin simulations with v-grad v in the $xy$, $xz$ and $yz$ planes for $Pe=100$ and $\beta=1$.
At this relatively high Péclet number the orientational distribution is strongly anisotropic. For the case with shear in the $xy$ plane the peak of the distribution is centered in the same plane. It does not point directly along the $x$ axis, but is rotated towards the $y$-axis. The spot is not symmetric, being somewhat stretched in the $z$ direction. It is clear that the remaining distributions $xz$ and $yz$ are related to the first one by a simple rotation. This helps to confirm the correct functioning of the Langevin algorithm as well as the graphical representation.
In Fig. 5 (left) we show the effect of varying the Péclet number between 1 and 100 for particles with shape parameters $\beta=0.5$, $0.8$ and $1$. As expected, with increasing noise (decreasing $Pe$) the distributions spread. For a fixed value of $Pe$ the distribution becomes more diffuse as the particle becomes less elongated.
These numerical solutions can be compared with the solutions of the Fokker-Planck equation. Fig. 5 (right) shows the probability surfaces for the solution of the Fokker-Planck equation (29) with $l_{max}=10$ for the same Péclet numbers. The solution was obtained with shear in the $xz$ plane.
Figure 5: Left: Density plots of the particle orientation calculated from Langevin simulations with v-grad v in the $xz$ plane, right: Probability surface of the particle orientation calculated from the solution of the Fokker-Planck equation, Eq. (29) with $l_{max}=10$. For the probability surfaces the distance of the surface from the origin in a given direction is proportional to the probability of the particle orientation in that direction.
@@ V.4 Order Matrix
Figure 8: Scalar order parameters $s$ (blue) and $r$ (green) of Jeffery orbits without noise averaged over the initial orientation as a function of $\beta$.
@@ VII.4 Orientational Order Matrix
Figure 14: (a) Angle for which the distribution displays a maximum and (b) $\theta_{p}$ as a function of the noise for $e=8,4,2,1.5$ top-to-bottom calculated from the exact solution. For large noise, both approach $\pi/4$ (dashed line).
@@ Appendix A From Stratonovich to Itô expression of the Langevin equation
$\frac{df\left(\textbf{p}\right)}{dt}=\bm{\nabla}f\cdot\textbf{A}+\bm{\nabla}f\cdot\textbf{B}\bm{\xi}+\frac{1}{2}\sum_{i,j}\left(BB^{\top}\right)_{ij}\partial_{i}\partial_{j}f\left(\textbf{p}\right)$ (77)
@@
where
$g_{m}=\left\{\begin{array}[]{ccc}\sqrt{2}&&\textrm{ if }m=0\\ 1&&\textrm{otherwise}\end{array}\right.$ (100)
@@ Appendix C Perturbation theory in 3D using $Pe$ as the expansion parameter
$\Lambda\psi_{0}$ $=$ $\nabla^{2}\psi_{1}$
$\Lambda\psi_{1}$ $=$ $\nabla^{2}\psi_{2}$ (106)
@@
which gives the scalar order parameter
@@
Note that for small $Pe$ (or large $D/\dot{\gamma}$) Eq. (123) reduces to this expression. Similarly we find
@@ Appendix D Perturbation theory in 3D using $\beta$ as the expansion parameter
Substituing in the Langevin equation gives
@@
$\mathbf{Q}=\begin{pmatrix}\frac{\beta}{5\sqrt{1+36\epsilon^{2}}}&0&\frac{6(\epsilon)\beta}{5\sqrt{1+36\epsilon^{2}}}\\ 0&0&0\\ \frac{6(\epsilon)\beta}{5\sqrt{1+36\epsilon^{2}}}&0&-\frac{\beta}{5\sqrt{1+36\epsilon^{2}}}\end{pmatrix}$ (121)
@@
Appendix E The scalar order parameters $s$ and $r$ for Jeffery orbits with a random initial orientation

### v2

# Exploring noisy Jeffery orbits: A combined Fokker-Planck and Langevin analysis in 2D and 3D

@@ Abstract
The behavior of non-spherical particles in a shear-flow is of significant practical and theoretical interest. These systems have been the object of numerous investigations since the pioneering work of Jeffery a century ago. His eponymous orbits describe the deterministic motion of an isolated, rod-like particle in a shear flow. Subsequently, the effect of adding noise was investigated. The theory has been applied to colloidal particles, macromolecules, anisometric granular particles and most recently to microswimmers, for example bacteria. We study the Jeffery orbits of elongated (uniaxial, prolate) particles subject to noise using Langevin simulations and a Fokker-Planck equation. We extend the analytical solution for infinitely thin needles ($\beta=1$) obtained by Doi and Edwards to particles with arbitrary shape factor ($0\leq\beta\leq 1$) and validate the theory by comparing it with simulations. We examine the rotation of the particle around the vorticity axis and study the orientational order matrix. We use the latter to obtain scalar order parameters $s$ and $r$ describing nematic ordering and biaxiality from the orientational distribution function. The value of $s$ (nematic ordering) increases monotonically with increasing Péclet number, while $r$ (measure of biaxiality) displays a maximum value. From perturbation theory we obtain simple expressions that provide accurate descriptions at low noise (or large Péclet numbers). We also examine the orientational distribution in the v-grad v plane and in the perpendicular direction. Finally we present the solution of the Fokker-Planck equation for a strictly two-dimensional (2D) system. For the same noise amplitude the average rotation speed of the particle in 3D is larger than in 2D.
@@ I Introduction
Even a century after its publication, Jeffery’s work is still relevant to many current research fields for a deep understanding of the behavior of sheared elongated particles, including rheology of granular flows and suspensions [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], nanoparticles, e.g. silica micro-rods in channels [15, 16, 17], red blood cells [18], as well as self-assembly and systems of active particles such as microswimmers [19, 20, 21, 22, 23, 24, 22, 23, 25, 26, 27, 28].
@@ II Jeffery Orbit
where ${\bf p}$ is a unit vector specifying the particle orientation, ${\bf E}=(\nabla{\bf u}+\nabla{\bf u}^{T})/2$, ${\bf W}=(\nabla{\bf u}-\nabla{\bf u}^{T})/2$ (with the convention $\left(\nabla{\bf u}\right)_{ij}=\partial_{j}u_{i}$) are the strain rate and vorticity tensors and $\beta$ is the shape factor (Bretherton parameter [29])
@@
for a particle of elongation $e$. In this article we restrict our attention to the uniaxial, prolate case, $\beta>0$ for which $e$ is the ratio of the longest axis to the shortest one. See Fig. 1. Note that $\beta=1$ and $\beta=0$ correspond to infinitely elongated and spherical particles respectively. Particular realizations of the Jeffery equations have been obtained by choosing the flow geometry. See, for example [1, 44, 26]. For shear in the $xy$ plane, i.e., ${\bf u}=\dot{\gamma}y\hat{\bf x}$, substituting in (1) and using standard spherical polar coordinates ($\theta,\;\phi$)
Figure 1: One of the coordinate systems used to describe the motion of an ellipsoid of revolution. In this case with shear in the $xy$ (v-grad v) plane. The $x,y,z$ axes correspond to the flow, gradient and vorticity directions, respectively. We also used other systems with shear in the $zz$ and $yz$ planes (not shown).
@@
$e\tan\phi=-\tan\left(\frac{\dot{\gamma}t}{e+1/e}+\kappa\right),$ (5)
@@ III Langevin Description
To simulate the Jeffery orbits with noise, these equations can be combined with those for the Jeffery orbits [19] (the Itô drift being unchanged). For shear in the $xy$ plane we obtain
@@ IV Fokker-Planck description
where $D$ is the rotational diffusion coefficient, assumed constant [42, 44], and where ${\bf w}=({\bf I}-{\bf p}\otimes{\bf p})(\beta{\bf E}+{\bf W}){\bf p}$, i.e., the deterministic part of the Langevin equation (10).
In spherical coordinates, ${\bf w}(\theta,\phi)=(0,\dot{\theta},\dot{\phi}\sin\theta)$ and
$\bm{\nabla}\cdot\left(\bf w\psi\right)=\frac{1}{\sin\theta}\partial_{\theta}\left(\dot{\theta}\sin\left(\theta\right)\psi\right)+\partial_{\phi}\left(\dot{\phi}\psi\right)$ (20)
It is particularly convenient to solve the Fokker-Planck equation with shear (v-grad v) in the $xz$ plane, which exploits the higher symmetry of the system with spherical coordinates (the reflection symmetry in the $xz$ plane allows one to use real spherical harmonics). Substituting (8) and (9) in (20), we obtain for the steady state
@@
$\hat{\Lambda}=-\frac{3}{2}\beta\sin(2\theta)\cos\phi\;+\left(\frac{1-\beta}{2}+\beta\cos^{2}\theta\right)\cos\phi\frac{\partial}{\partial\theta}-\frac{1+\beta}{2}\cot\theta\sin\phi\frac{\partial}{\partial\phi}$ (22)
@@
$\hat{\Lambda}=\beta\hat{\Lambda}_{1}+\frac{1-\beta}{2}i\hat{L}_{y},$ (25)
a result that, to the best of our knowledge, has not been given previously. For $\beta=0$, corresponding to a sphere, Eq. (21) simplifies to
@@
$-\epsilon\,l(l+1)b_{lm}=\sum_{l^{\prime}\geq 0,\;{\rm even}}^{l_{max}}\sum_{m^{\prime}=0}^{l^{\prime}}(lm|\hat{\Lambda}|l^{\prime}m^{\prime})b_{l^{\prime}m^{\prime}}$ (31)
for $l\leq l_{max}$. From this we generate $l_{max}(1+l_{max}/4)$ simultaneous equations that can be solved to find the coefficients $\{b_{lm}\}$.
@@ V.2 Rotation around the vorticity axis
We first examine the rotation of the particle around the vorticity axis. For shear in the $xy$ plane, in the steady state we have $\langle d\theta/dt\rangle=0$, where the angular brackets denote a time average, giving
@@
With zero noise, expressing $\phi$ as a function of time using Eq. (5) and integrating over one period we obtain
@@
The rate of rotation increases with decreasing elongation reaching a maximum value of $\dot{\gamma}/2$ for spherical particles, $\beta=0$. Infinitely thin particles, $\beta=1$ are singular in that they do not rotate periodically. Indeed $\phi=0$ is an attractor of the differential equation (4): the particle eventually aligns with the flow for all starting orientations.
The effect of noise on the rate of rotation is shown in Fig. 3. As in the noiseless case, for a fixed level of noise the rate of rotation increases with decreasing elongation and, for fixed elongation ($\beta$), the rate of rotation increases with increasing noise. Perturbation theory, as developed in Appendix C, can be used to calculate the rotation speed in the limit of large noise (or small Pe).
@@ V.3 Orientational distribution function
Figure 4: Density plots of the particle orientation from Langevin simulations with v-grad v in the $xy$, $xz$ and $yz$ planes for $Pe=100$ and $\beta=1$. The contours show the solution of the Fokker-Planck equation (30) with $l_{max}=16$.
At this relatively high Péclet number the orientational distribution is strongly anisotropic. For the case with shear in the $xy$ plane the peak of the distribution is centered in the same plane. It does not point directly along the $x$ axis, but is rotated towards the $y$-axis [13]. The spot is not symmetric, being somewhat stretched in the $z$ direction. It is clear that the remaining distributions $xz$ and $yz$ are related to the first one by a simple rotation. This helps to confirm the correct functioning of the Langevin algorithm as well as the graphical representation.
In Fig. 5 (left) we show the effect of varying the Péclet number between 1 and 100 for particles with shape parameters $\beta=0.5$, $0.8$ and $1$. As expected, with increasing noise (decreasing $Pe$) the distributions spread. For a fixed value of $Pe$, the distribution becomes more diffuse as the particle becomes less elongated. We also show the solutions of the Fokker-Planck equation (30) as contours superimposed on the density plots. Fig. 5 (right) shows the same solutions as probability surfaces. The solution was obtained with shear in the $xz$ plane. The full symmetry of the distributions is more clearly apparent in this probability surface representation.
Figure 5: Left: Density plots of the particle orientation calculated from Langevin simulations with v-grad v in the $xz$ plane. The contours show the solution of the Fokker-Planck equation (30) with $l_{max}=16$, right: The same Fokker-Planck solutions shown as probability surfaces. The distance of the surface from the origin in a given direction is proportional to the probability of the particle orientation in that direction.
@@ V.4 Order Matrix
Figure 8: Scalar order parameters $s$ (blue) and $r$ (green) of Jeffery orbits without noise averaged over the initial (isotropic) orientation as a function of $\beta$.
@@ VII.4 Orientational Order Matrix
Figure 14: Angle for which the distribution displays a maximum $\theta_{\rm max}$, (blue, solid) and $\theta_{p}$ (black, dashed) as a function of the noise for $e=8,4,2,1.5$ top-to-bottom calculated from the exact solution. For large noise, both approach $\pi/4$ (dashed line).
@@ Appendix A From Stratonovich to Itô expression of the Langevin equation
$\frac{df\left(\textbf{p}\right)}{dt}=\bm{\nabla}f\cdot\textbf{A}+\bm{\nabla}f\cdot\textbf{B}\bm{\xi}+\frac{1}{2}\sum_{i,j}\left(\textbf{B}\textbf{B}^{\top}\right)_{ij}\partial_{i}\partial_{j}f\left(\textbf{p}\right)$ (78)
@@
where, as previously stated
$g_{m}=\left\{\begin{array}[]{ccc}\sqrt{2}&&\textrm{ if }m=0\\ 1&&\textrm{ if }m>0\\ 0&&\textrm{ if }m<0\end{array}\right.$ (101)
@@ Appendix C Perturbation theory in 3D using $Pe$ as the expansion parameter
$\hat{\Lambda}\psi_{0}$ $=$ $\nabla^{2}\psi_{1}$
$\hat{\Lambda}\psi_{1}$ $=$ $\nabla^{2}\psi_{2}$ (107)
@@
which gives the scalar order parameter (defined in section D)
@@
Similarly we find
@@ Appendix D Perturbation theory in 3D using $\beta$ as the expansion parameter
Substituting in the Fokker-Planck equation gives
@@
$\mathbf{Q}=\begin{pmatrix}\frac{\beta}{5(1+36\epsilon^{2})}&0&\frac{6(\epsilon)\beta}{5(1+36\epsilon^{2})}\\ 0&0&0\\ \frac{6(\epsilon)\beta}{5(1+36\epsilon^{2})}&0&-\frac{\beta}{5(1+36\epsilon^{2})}\end{pmatrix}$ (122)
@@
Note that for small $Pe$ (or large $D/\dot{\gamma}$) Eq. (111) reduces to this expression.
Appendix E The scalar order parameters $s$ and $r$ for Jeffery orbits with a uniformly distributed random initial orientation